OpenStudy (anonymous):

Simplify the following expression. 41 - (-26) A. 67 B. -15 C. 15 D. -67 @jabez177

OpenStudy (jojo12g):

it a 41−(−26) =41−(−26) =41+26 =67

OpenStudy (jojo12g):

do u under stand

OpenStudy (jojo12g):

41−(−26) =41−(−26) =41+26 =67

OpenStudy (lolacole12):

a

OpenStudy (lolacole12):

41−(−26) =41−(−26) =41+26 =67

OpenStudy (anonymous):

-----8 Count the number of - signs. If the total is an even number then the sign of 8 is +, otherwise -. -----8 = -8

OpenStudy (anonymous):

thz

OpenStudy (anonymous):

i got it thx